OpenStudy (anonymous):

Hey guys so im stuck on a problem!! it is x^2-5x+1=0 so far i got x^2-5x+25/4=-1(25/4) what i want to know is how to Write the left side as the square of a binomial. Hey guys so im stuck on a problem!! it is x^2-5x+1=0 so far i got x^2-5x+25/4=-1(25/4) what i want to know is how to Write the left side as the square of a binomial. @Mathematics

OpenStudy (asnaseer):

\[(x-5/2)^2=x^2-5x+25/4\]

OpenStudy (anonymous):

btw the right side is -1 +25/4 makes a HUGE DIFFERENCE

OpenStudy (asnaseer):

correct - I was answering your "left-side" question.

OpenStudy (anonymous):

oh okay thanks!

OpenStudy (asnaseer):

:-) yw

OpenStudy (anonymous):

MEtal for you (^.^)

OpenStudy (radar):

What next?

OpenStudy (anonymous):

hummm? i got the rest

OpenStudy (radar):

(x-5/2)^2=21/4 Good, keep up the good work

OpenStudy (anonymous):

thanks